# Magnetic helicity and energy budgets of jet events from an emerging solar active region

### v1

@@ 4 Conclusions
In this paper we used the CB method (Georgoulis et al., 2012) to evaluate the free magnetic energy and helicity budgets of a small emerging AR. The AR was bipolar and during its emergence phase which lasted 47 hours produced no flares above C1.0-class or CMEs. However, we were able to identify by visual inspection of 211-Å movies 60 jets emanating from the AR.
@@
Throughout the interval we studied, the $E_{f}$ and $H$ budgets of the AR were below established thersholds (see Tziotziou et al., 2012; Liokati et al., 2022) which, if crossed, the AR is likely to erupt. The time profiles of the free magnetic energy and helicity resulted from the superposition of two components: (i) A slowly-varying one which was broadly consistent with the evolution of both the total unsigned magnetic flux of the AR and the connected magnetic flux. (ii) Discrete localized peaks of much shorter duration (full widths at half maximum of $\sim$55-160 min). Eight such peaks in each of the $E_{f}$ and $H$ curves (all well beyond uncertainties) occurred co-temporally with jet events produced in the AR. These local helicity peaks can reasonably be attributed to the jet events because no other type of eruptive activity was registered in the AR. Furthermore, their pairing with $E_{f}$ localized peaks supports the same conclusion for the origin of the simultaneous $E_{f}$ peaks.
@@
The jets associated with localized peaks in the $E_{f}$ and $H$ budgets of the AR are distinguished from the other AR jets by their larger areas of their bases as well as by their longer durations. The former is in line with the fact that $E_{f}$ and $H$ are extensive quantities. It is interesting, though, that these major jets are also associated with local peaks in the time profiles of the normalized free magnetic energy and helicity parameters.
@@
The free magnetic energy and helicity losses associated with the jets are in the ranges of $(1.1-6.9)\times 10^{29}$ erg and $(0.5-7.1)\times 10^{40}$ Mx2, respectively. These values are one to two orders of magnitude smaller than the relevant changes associated with CMEs (see Liokati et al., 2023, and references therein). The derived free energy losses are consistent with the high-end limits of the thermal energy of jets (see Shen, 2021). There are no previous explicit reports based on observations about helicity changes directly associated with jets. The percentage losses associated with the jets are significant: 9-57% for the normalized free magnetic energy and 13-76% for the normalized helicity. There is a trend the percentage losses to be larger early on in the evolution of the AR (see Table 1).
@@
This is the first report where changes in the magnetic free energy and helicity budgets of an AR are registered with such clarity with jet activity. Green et al. (2022) were the first to report jet activity in intervals when the helicity index (see Sect. 1) attains large values. In our study the close synchronization of localized $E_{f}$ and $H$ peaks with jet events allowed us to estimate, for the first time, the free magnetic energy and helicity budgets associated with individual jet events.
@@
More case studies are required to check how often jets may have significant imprints in the evolution of the free magnetic energy and helicity budgets of emerging ARs. As more case studies will be accumulated, it would be interesting to investigate whether the collective $E_{f}$ and $H$ budgets of such jets are significant. Computations of the field line helicity (e.g. Yeates & Page, 2018; Moraitis et al., 2019, 2021, 2024) may also provide insights about the distribution of helicity over the different components of individual jets.
@@ Appendix A Jets produced in AR11096
In panel (b) of Fig. A.2 we show the duration of the jets. The duration is defined as the interval between the last and first appearance of the jet spire in the AIA images. The duration of jets appearing in only one AIA image has been somewhat arbitrarily set to 2 min (i.e. the cadence of our AIA datacube).

### v2

@@ 4 Conclusions
In this work we used the CB method (Georgoulis et al., 2012) to evaluate the free magnetic energy and helicity budgets of a small emerging AR. The AR was bipolar and during its emergence phase (which lasted 47 hours) produced no flares above C1.0-class or CMEs. However, we were able to identify 60 jets emanating from the AR by visually inspecting 211-Å movies.
@@
Throughout the interval we studied, the $E_{f}$ and $H$ budgets of the AR were below established thresholds (see Tziotziou et al., 2012; Liokati et al., 2022); if these are crossed, the AR is likely to erupt. The time profiles of the free magnetic energy and helicity resulted from the superposition of two components: (i) a slowly varying one that was broadly consistent with the evolution of both the total unsigned magnetic flux of the AR and the connected magnetic flux; and (ii) discrete localized peaks of much shorter durations (full widths at half maximum of $\sim$55-160 min). Eight such peaks in each of the $E_{f}$ and $H$ curves (all well beyond uncertainties) occurred co-temporally with jet events produced in the AR. These local helicity peaks can reasonably be attributed to the jet events because no other type of eruptive activity was registered in the AR. Furthermore, their pairing with $E_{f}$ localized peaks supports the same conclusion for the origin of the simultaneous $E_{f}$ peaks.
@@
The jets associated with localized peaks in the $E_{f}$ and $H$ budgets of the AR are distinguished from the other AR jets by the larger areas of their bases and their longer durations. The former is in line with the fact that $E_{f}$ and $H$ are extensive quantities. It is interesting, though, that these major jets are also associated with local peaks in the time profiles of the normalized free magnetic energy and helicity parameters.
@@
The free magnetic energy and helicity losses associated with the jets are in the ranges of $(1.1-6.9)\times 10^{29}$ erg and $(0.5-7.1)\times 10^{40}$ Mx2, respectively. These values are one to two orders of magnitude smaller than the relevant changes associated with CMEs (see Liokati et al., 2023, and references therein). The derived free energy losses are consistent with the high-end limits of the thermal energy of jets (see Shen, 2021). There are no previous explicit reports based on observations about helicity changes directly associated with jets. The percentage losses associated with the jets are significant: 9-57% for the normalized free magnetic energy and 13-76% for the normalized helicity. There is a trend for the percentage losses to be larger early on in the evolution of the AR (see Table 1).
@@
This is the first report where changes in the magnetic free energy and helicity budgets of an AR are registered with such clarity with jet activity. Green et al. (2022) was the first to report jet activity in intervals when the helicity index (see Sect. 1) attains large values. In our study, the close synchronization of localized $E_{f}$ and $H$ peaks with jet events allowed us to estimate, for the first time, the free magnetic energy and helicity budgets associated with individual jet events.
@@
More case studies are required to check how often jets may have significant imprints in the evolution of the free magnetic energy and helicity budgets of emerging ARs. As more case studies accumulate, it will be interesting to investigate whether the collective $E_{f}$ and $H$ budgets of such jets are significant. Computations of the field line helicity (e.g. Yeates & Page, 2018; Moraitis et al., 2019, 2021, 2024) may also provide insights into the distribution of helicity over the different components of individual jets.
@@ Appendix A Jets produced in AR11096
In panel (b) of Fig. A.2 we show the duration of the jets. The duration is defined as the interval between the last and first appearance of the jet spire in the AIA images. The duration of jets appearing in only one AIA image has been somewhat arbitrarily set to 2 min (i.e., the cadence of our AIA datacube).